# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 1 Introduction
Large Language Models (LLMs), mainly characterized by their vast number of parameters, have emerged as a promising cornerstone for the development of natural language processing(NLP) and artificial intelligence (AI) Zhang et al. (2023). These models have demonstrated remarkable capabilities across a wide range of NLP applications. Foundation models like GPT-4 OpenAI et al. (2024), GPT-3 Brown et al. (2020), Qwen 2.5-VL Bai et al. (2025a), Deepseek-V3 DeepSeek-AI et al. (2025), Meta (LLaMA-3, LLaMA-2 Touvron et al. (2023b), LLaMA-1 Touvron et al. (2023a)), Gemini DeepMind (2023) are pre-trained on massive datasets and serve as the backbone for various AI tasks, from text generation to multimodal understanding—the ability to process and interpret multiple types of data, such as text and images. Their extensive factual knowledge has significantly improved their effectiveness in information retrieval, enabling them to provide more accurate and contextually relevant responses AlKhamissi et al. (2022); Petroni et al. (2019). As a result, these models achieve high accuracy on benchmark evaluations, demonstrating their strong predictive capabilities across diverse tasks OpenAI et al. (2024); Bai et al. (2025a); DeepSeek-AI et al. (2025); Grattafiori et al. (2024). However, accuracy alone is insufficient to ensure reliability in real-world applications.
@@
In recent years, despite improvements in the accuracy of deep neural network models, researchers have found that they are easy to fool by applying a specific imperceptible perturbation Szegedy et al. (2014). A deep learning model’s accuracy describes how accurately the model predicts sample points over a distribution. If a model cannot predict accurately, then no matter what properties this model holds, it would be meaningless. Although accuracy is the most fundamental metric in the evaluation and is essential to validate the performance of an LLM before it is released, it is equally important to test its robustness when deploying for the development of reliable AI systems. Robustness, a crucial aspect of deep learning models, describes their ability to maintain stable predictions when faced with specific perturbations or variations in the input data Liu et al. (2024b). Despite recent advancements, ensuring the robustness of LLMs remains a critical challenge. LLMs may encounter unpredictable variations in language, shifts in data (changes in the data patterns that a model encounters, which may affect its performance), and adversarial inputs that can affect their performance. In practical applications, these models often face noisy, unstructured text, breaking the assumption that input data will always be clean and well-formed. Additionally, a lack of robustness may result in unintended biases, incorrect predictions, or overreliance on spurious correlations (misleading correlations in data that a model may rely on incorrectly), raising concerns about their reliability in high-stakes environments such as healthcare Wan et al. (2024b; a); He et al. (2025), law Lee (2023); Cui et al. (2024); Yue et al. (2023); Bhambhoria et al. (2024); Bai et al. (2021b), complex reasoning Yang et al. (2023), code generation Sun et al. (2024b); Lin et al. (2024a); Mishra et al. (2024); Hassid et al. (2024), and finance Lee et al. (2025); Zhao et al. (2024a); Li et al. (2024c); Wu et al. (2023). Because of this, a thorough and exacting assessment of LLM robustness that goes beyond conventional accuracy criteria is required.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=6em, edge+=darkgray, line width=1pt, s sep=6pt, inner xsep=3pt, inner ysep=4pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em,font=, text centered, where level=2text width=12em,font=, text centered, where level=3text width=13em,font=, where level=4text width=33em,font=, align=left, [ A Survey on Robustness of LLMs, fill=mygrey!10, ver [ Robustness in
LLMs (§ 2), fill=blue!10 [ What is Robustness
in LLMs (§ 2.1), fill=blue!10 [ Guo et al. (2023); Du et al. (2022); Singh et al. (2024); Liu et al. (2024b); Mahaut et al. (2024)
Sarker et al. (2024); Ailem et al. (2024); Aljohani et al. (2025); Li et al. (2024d), align=left, text width=44.5em ] ] [ Key dimensions
of Robustness(§ 2.2), fill=blue!10 [ Chacko et al. (2024); Yang et al. (2024d); Tao et al. (2024); Gan et al. (2024); Xhonneux et al. (2024)
Jiang et al. (2025); Du et al. (2022); Yuan et al. (2023); Nagarajan et al. (2024); Luo et al. (2024)
Zhou et al. (2024b); Yang et al. (2025b); Lin et al. (2024b); Agrawal et al. (2025), align=left, text width=44.5em ] ] ] [ Sources of non-
robustness (§ 3), fill=yellow!10 [ Data-Related (§ 3.1), fill=yellow!10 [ Du et al. (2022); Li et al. (2024b); Jung et al. (2025); Ailem et al. (2024); Aljohani et al. (2025)
Mahaut et al. (2024); Liu et al. (2024b); Yan et al. (2024a); Singh et al. (2024); Zhou et al. (2024d), align=left, text width=44.5em ] ] [ Training-Related (§ 3.2), fill=yellow!10 [ Nagarajan et al. (2024); Yan et al. (2024b); Barnhart et al. (2025); Huang et al. (2024a)
Yuan et al. (2023); Chen et al. (2020); Yang et al. (2024b); Lin et al. (2024b), align=left, text width=44.5em ] ] [ Architectural (§ 3.3), fill=yellow!10 [ Zhang (2023); Das et al. (2025); Ma et al. (2025); Fan & Tao (2024); Anwar et al. (2024), align=left, text width=44.5em ] ] [ Inference-Related
(§ 3.4), fill=yellow!10 [ Nik et al. (2025); Bang et al. (2023); Shi et al. (2024a); Yao et al. (2024); Shen et al. (2024)
Yu et al. (2024), align=left, text width=44.5em ] ] ] [ Mitigation Strategies
(§ 4), fill=orange!10 [ Pre-processing
Approaches (§ 4.1), fill=orange!10 [ Data Augmentation, fill=orange!10, text centered [ Dong et al. (2021); Weng (2023); Zhao et al. (2020); Jiang et al. (2020)
Xu et al. (2023); Liu & Sun (2023); Lee et al. (2024b)
Qian et al. (2022); Zayed et al. (2022); Lim et al. (2023) ] ] [ Data Filtering, fill=orange!10, text centered [ Huang et al. (2024a); Laurençon et al. (2023); Dong et al. (2024a)
Li et al. (2025b); Bendinelli et al. (2025); Chen et al. (2023a) ] ] ] [ In-processing
Approaches (§ 4.2), fill=orange!10 [ Adversarial Training, fill=orange!10, align=center [ Bai et al. (2021a); Liu & Sun (2023); Altinisik et al. (2023)
Xhonneux et al. (2024); Dong et al. (2021); Rafailov et al. (2024)
Rahimian & Mehrotra (2022); Kuhn et al. (2025); Wu et al. (2025) ] ] [ Regularization, fill=orange!10, text centered [ Jiang et al. (2020); Xu et al. (2023); Mahabadi et al. (2020)
Utama et al. (2020); Garimella et al. (2021); Xu et al. (2025)
Ravichandran et al. (2020); Li et al. (2022) ] ] [ Alignment, fill=orange!10, text centered [ Oh & Demberg (2025); Ouyang et al. (2022); Christiano et al. (2023)
Hong et al. (2024); Ethayarajh et al. (2024); Gupta et al. (2025)
Tran et al. (2025); Sun (2024); Herrera-Poyatos et al. (2025)
Zhong et al. (2025a); Xu et al. (2023); Wang et al. (2024a)
Dong et al. (2024b); Xiong et al. (2024); Zhao et al. (2024b) ] ] ] [ Intra-processing
Approaches (§ 4.3), fill=orange!10 [ Robust Prompting, fill=orange!10, text centered [ Agrawal et al. (2025); Chen et al. (2025b); Li et al. (2024d) ] ] [ Weight
Redistribution, fill=orange!10, text centered [ Chai & Wang (2022); Zhong et al. (2025b) ] ] [ Modified Decoding, fill=orange!10, text centered [ Beyer et al. (2025); Kumar et al. (2025a); Gu et al. (2024)
Daheim et al. (2025) ] ] [ Inference Time
Adaptation, fill=orange!10, text centered [ Manvi et al. (2024); Snell et al. (2024); Lad et al. (2024)
Jha & Reagen (2024); Ye et al. (2025) ] ] ] [ Post-processing
Approaches (§ 4.4), fill=orange!10 [ Output Filtering, fill=orange!10, text centered [ Kumar et al. (2025b); Liu et al. (2023); Phute et al. (2024)
Weng (2023); Rebedea et al. (2023); Wang et al. (2025) ] ] [ LLM-as-a-judge, fill=orange!10, text centered [ Gu et al. (2025); Kumar et al. (2025b); Cantini et al. (2025)
Raina et al. (2024); Beyer et al. (2025); Li et al. (2025c) ] ] ] ] [ Metrics and
Benchmarks (§ 5), fill=orange!10 [ Metric, fill=orange!10, text centered ] [ Benchmarks, fill=orange!10, text centered ] ] [ Future Work (§ 7), fill=green!10 [ Will be updated later, text width=36em, fill=green!10 ] ] ]
Figure 1: The outline of the survey on robustness of large language models.
@@
However, advancing research on the robustness of LLMs also faces significant challenges. First, the absence of standardized definitions and categorizations of robustness complicates cross-study comparisons, as research efforts often differ in how they frame their scope and components. Second, pinpointing the root causes of non-robustness, such as data biases, unstable training dynamics, architectural limitations, or inference inefficiencies, requires disentangling complex interactions across these factors. Third, while numerous mitigation strategies exist, their effectiveness often depends on addressing specific vulnerabilities, making it difficult to generalize solutions. Fourth, evaluating robustness remains fragmented, with no unified metrics or benchmarks to holistically assess performance under diverse real-world conditions like noisy inputs or adversarial attacks. Finally, existing literature on LLM robustness (Zhao et al., 2025; Gu et al., 2025; He et al., 2025; Liu et al., 2024b; Mehrabi et al., 2022; Wang et al., 2025; Lu et al., 2024; Weng, 2023; Yang et al., 2024a) is scattered across domains, limiting cross-pollination of insights and hindering the development of cohesive solutions.
@@ 2 Robustness in LLMs
In the context of LLMs from various literatures, robustness refers to their ability to perform reliably and safely despite challenges like input variations, unexpected data, or adversarial attacks. It is frequently discussed as a critical component within the broader framework of AI Trustworthiness. Trustworthy AI systems are expected to be reliable, safe, fair, transparent, accountable, and robust. Robustness underpins several of these qualities; for instance, a model prone to adversarial manipulation (lack of robustness) cannot be considered safe or reliable (Wang et al., 2023; Aljohani et al., 2025). Similarly, a model that hallucinates or provides inconsistent answers under slightly different prompts lacks reliability, another facet closely tied to robustness (Zhou et al., 2024b). Research into the vulnerabilities of machine learning models, including LLMs, has often focused on adversarial robustness - the model’s resilience against inputs intentionally crafted to cause misclassification or undesired behaviour (Zhao et al., 2025).

### v2

@@ 1 Introduction
Large Language Models (LLMs), mainly characterized by their vast number of parameters, have emerged as a promising cornerstone for the development of natural language processing(NLP) and artificial intelligence (AI) unknown. These models have demonstrated remarkable capabilities across a wide range of NLP applications. Foundation models like GPT-4 openai2024gpt4technicalreport, GPT-3 brown2020languagemodelsfewshotlearners, Qwen 2.5-VL bai2025qwen25vltechnicalreport, Deepseek-V3 deepseekai2025deepseekv3technicalreport, Meta (LLaMA-3, LLaMA-2 touvron2023llama2openfoundation, LLaMA-1 Touvron2023LLaMAOA), Gemini gemini1 are pre-trained on massive datasets and serve as the backbone for various AI tasks, from text generation to multimodal understanding—the ability to process and interpret multiple types of data, such as text and images. Their extensive factual knowledge has significantly improved their effectiveness in information retrieval, enabling them to provide more accurate and contextually relevant responses alkhamissi2022reviewlanguagemodelsknowledge; petroni2019languagemodelsknowledgebases. As a result, these models achieve high accuracy on benchmark evaluations, demonstrating their strong predictive capabilities across diverse tasks openai2024gpt4technicalreport; bai2025qwen25vltechnicalreport; deepseekai2025deepseekv3technicalreport; grattafiori2024llama3herdmodels. However, accuracy alone is insufficient to ensure reliability in real-world applications.
@@
In recent years, despite improvements in the accuracy of deep neural network models, researchers have found that they are easy to fool by applying a specific imperceptible perturbation szegedy2014intriguingpropertiesneuralnetworks. A deep learning model’s accuracy describes how accurately the model predicts sample points over a distribution. If a model cannot predict accurately, then no matter what properties this model holds, it would be meaningless. Although accuracy is the most fundamental metric in the evaluation and is essential to validate the performance of an LLM before it is released, it is equally important to test its robustness when deploying for the development of reliable AI systems. Robustness, a crucial aspect of deep learning models, describes their ability to maintain stable predictions when faced with specific perturbations or variations in the input data liu2024trustworthyllmssurveyguideline. Despite recent advancements, ensuring the robustness of LLMs remains a critical challenge. LLMs may encounter unpredictable variations in language, shifts in data (changes in the data patterns that a model encounters, which may affect its performance), and adversarial inputs that can affect their performance. In practical applications, these models often face noisy, unstructured text, breaking the assumption that input data will always be clean and well-formed. Additionally, a lack of robustness may result in unintended biases, incorrect predictions, or overreliance on spurious correlations (misleading correlations in data that a model may rely on incorrectly), raising concerns about their reliability in high-stakes environments such as healthcare wan2024gmapgeneralmemoryaugmentedpretrained; wan2024medunicunifyingcrosslingualmedical; he2025surveylargelanguagemodels, law lee2023lexgpt01pretrainedgptj; cui2024chatlawmultiagentcollaborativelegal; yue2023disclawllmfinetuninglargelanguage; bhambhoria2024evaluatingailawbridging; bai2021societal, complex reasoning yang2023harnessingpowerllmspractice, code generation sun2024enhancingcodegenerationperformance; lin2024soen101codegenerationemulating; mishra2024granitecodemodelsfamily; hassid2024largerbetterimprovedllm, and finance Lee_2025; zhao2024revolutionizingfinancellmsoverview; li2024largelanguagemodelsfinance; wu2023bloomberggptlargelanguagemodel. Because of this, a thorough and exacting assessment of LLM robustness that goes beyond conventional accuracy criteria is required.
[…]
A Survey on Robustness of LLMs
Robustness in LLMs (§ 2) What is Robustness in LLMs (§ 2.1) guo2023evaluatinglargelanguagemodels; du2022robustness; singh2024robustnessllmsperturbationstext; liu2024trustworthyllmssurveyguideline; Mahaut_2024 sarker2024syntacticrobustnessllmbasedcode; ailem2024examiningrobustnessllmevaluation; aljohani2025comprehensivesurveytrustworthinesslarge; li-etal-2024-evaluating-instruction Key dimensions of Robustness(§ 2.2) chacko2024adversarialattackslargelanguage; yang2024assessingadversarialrobustnesslarge; tao2024robustnesslargelanguagemodels; gan2024reasoningrobustnessllmsadversarial; xhonneux2024efficientadversarialtrainingllms jiang2025surveyadversarialrobustnessmultimodal; du2022robustness; yuan2023revisiting; nagarajan2024understandingfailuremodesoutofdistribution; luo2024robustftrobustsupervisedfinetuning zhou2024trustworthinessretrievalaugmentedgenerationsystems; yang2025enhancingsemanticconsistencylarge; lin2024mitigatingalignmenttaxrlhf; agrawal2025enhancingllmrobustnessperturbed Sources of non- robustness (§ 3) Data-Related (§ 3.1) du2022robustness; li-etal-2024-gsm; jung2025flexbenchmarkevaluatingrobustness; ailem2024examiningrobustnessllmevaluation; aljohani2025comprehensivesurveytrustworthinesslarge Mahaut_2024; liu2024trustworthyllmssurveyguideline; yan2024contrastiveinstructiontuning; singh2024robustnessllmsperturbationstext; zhou2024languagemodelsperformrobust Training-Related (§ 3.2) nagarajan2024understandingfailuremodesoutofdistribution; yan2024rewardrobustrlhfllms; barnhart2025aligningwhatlimitsrlhf; huang2024trustllmtrustworthinesslargelanguage yuan2023revisiting; chen2020distillingknowledgelearnedbert; yang2024confidencecalibrationrationalizationllms; lin2024mitigatingalignmenttaxrlhf Architectural (§ 3.3) zhang2023awesome; das2025genbfaevolutionaryoptimizationapproach; ma2025modelhemorrhagerobustnesslimits; fan2024resilientefficientllmscomparative; anwar2024adversarialrobustnessincontextlearning Inference-Related (§ 3.4) nik2025energyconsciousllmdecodingimpact; bang2023multitaskmultilingualmultimodalevaluation; shi2024thoroughexaminationdecodingmethods; yao2024retroformerretrospectivelargelanguage; shen-etal-2024-assessing yu2024rankragunifyingcontextranking Mitigation Strategies (§ 4) Pre-processing Approaches (§ 4.1) Data Augmentation dong-etal-2021-data; weng2023attack; Zhao_Liu_Peng_Metaxas_2020; jiang2020robustpretrainingadversarialcontrastive xu2023efficientadversarialcontrastivelearning; liu-sun-2023-end; Lee2024HarmAugED qian2022perturbationaugmentationfairernlp; zayed2022deeplearninghealthydata; 10203320 Data Filtering huang2024trustllmtrustworthinesslargelanguage; laurençon2023bigsciencerootscorpus16tb; dong2024baichuanseedsharingpotentialextensive li2025trustworthymachinelearningmemorization; bendinelli2025exploringllmagentscleaning; chen2023datajuiceronestopdataprocessing In-processing Approaches (§ 4.2) Adversarial Training bai2021recentadvancesadversarialtraining; liu-sun-2023-end; altinisik-etal-2023-impact xhonneux2024efficientadversarialtrainingllms; dong-etal-2021-data; rafailov2024directpreferenceoptimizationlanguage Rahimian_2022; kuhn2025distributionallyrobustoptimization; wu2025towards Regularization jiang2020robustpretrainingadversarialcontrastive; xu2023efficientadversarialcontrastivelearning; mahabadi2020endtoendbiasmitigationmodelling utama2020mindtradeoffdebiasingnlu; Garimella2021HeIV; xu2025collapsedlanguagemodelspromote ravichandran2020fairxgboostfairnessawareclassificationxgboost; li2022accuratefairnessimprovingindividual Alignment Oh_Demberg_2025; ouyang2022traininglanguagemodelsfollow; christiano2023deepreinforcementlearninghuman hong2024orpomonolithicpreferenceoptimization; ethayarajh2024ktomodelalignmentprospect; gupta2025alphaporewardshape tran2025vulnerabilitymitigationsafetyalignedlanguage; sun2024supervisedfinetuninginversereinforcement; herrerapoyatos2025overviewmodeluncertaintyvariability zhong2025comprehensivesurveyrewardmodels; xu2023efficientadversarialcontrastivelearning; ArmoRM dong2024rlhf; xiong2024iterative; zhao-etal-2024-improving Intra-processing Approaches (§ 4.3) Robust Prompting agrawal2025enhancingllmrobustnessperturbed; chen2025robustnessreferencingdefendingprompt; li-etal-2024-evaluating-instruction Weight Redistribution pmlr-v162-chai22a; zhong2025optimizingrlhftraininglarge Modified Decoding beyer2025llmsafetyevaluationslackrobustness; kumar2025certifyingllmsafetyadversarial; gu2024charedcharacterwiseensembledecoding daheim2025uncertaintyaware Inference Time Adaptation manvi2024adaptiveinferencetimecomputellms; snell2024scalingllmtesttimecompute; lad2024remarkablerobustnessllmsstages jha2024aerosoftmaxonlyllmsefficient; ye2025flashinferefficientcustomizableattention Post-processing Approaches (§ 4.4) Output Filtering kumar2025llmposttrainingdeepdive; liu2023promptingframeworkslargelanguage; phute2024llmselfdefenseself weng2023attack; rebedea2023nemoguardrailstoolkitcontrollable; wang2025surveyresponsiblellmsinherent LLM-as-a-judge gu2025surveyllmasajudge; kumar2025llmposttrainingdeepdive; cantini2025benchmarkingadversarialrobustnessbias raina-etal-2024-llm; beyer2025llmsafetyevaluationslackrobustness; li2025wantedknowllmbasedvulnerability Metrics and Benchmarks (§ 5) Metric Benchmarks Future Work (§ 6) Scalability of Defences, Robustness-Utility Trade-offs, Evaluation Gaps, Theoretical Understanding, and Real-world Deployment
Figure 1: The outline of the survey on robustness of LLMs.
@@ A Survey on Robustness of LLMs
However, advancing research on the robustness of LLMs also faces significant challenges. First, the absence of standardized definitions and categorizations of robustness complicates cross-study comparisons, as research efforts often differ in how they frame their scope and components. Second, pinpointing the root causes of non-robustness, such as data biases, unstable training dynamics, architectural limitations, or inference inefficiencies, requires disentangling complex interactions across these factors. Third, while numerous mitigation strategies exist, their effectiveness often depends on addressing specific vulnerabilities, making it difficult to generalize solutions. Fourth, evaluating robustness remains fragmented, with no unified metrics or benchmarks to holistically assess performance under diverse real-world conditions like noisy inputs or adversarial attacks. Finally, existing literature on LLM robustness (zhao2025surveylargelanguagemodels; gu2025surveyllmasajudge; he2025surveylargelanguagemodels; liu2024trustworthyllmssurveyguideline; mehrabi2022surveybiasfairnessmachine; wang2025surveyresponsiblellmsinherent; lu2024recentadvancesooddetection; weng2023attack; yang2024generalizedoutofdistributiondetectionsurvey) is scattered across domains, limiting cross-pollination of insights and hindering the development of cohesive solutions.
@@
The existence of numerous surveys, while indicating intense research activity, also suggests a potentially fragmented landscape where different communities might use varying terminology or focus on specific aspects of the broader robustness problem. A key contribution of a comprehensive survey like this one is to bridge these perspectives and synthesize the common underlying principles and challenges. Furthermore, existing surveys often highlight gaps and open questions that remain pertinent. The lack of unified definitions and practical frameworks for operationalising trustworthiness (including robustness) is noted (decerqueira2025mappingtrustworthinesslargelanguage). The difficulty in evaluating generative tasks robustly is frequently mentioned. Understanding and mitigating more subtle or complex shortcuts beyond simple lexical cues remains an area for research (howe2025scalingtrendslanguagemodel; zhou2024navigatingshortcutmazecomprehensive; yuan2023revisiting). Finally, ensuring that robustness improvements scale effectively with increasing model size is an ongoing concern.
1.1 Search Methodology
The purpose of this survey is to gather and categorize research work that helps to improve robustness in LLMs. A systematic search was conducted across multiple digital libraries and repositories to identify relevant studies on robustness in LLMs as shown in Fig. 2. Table 1 summarizes the databases and sources that were included:
Table 1: Search results by source
Source Identified Screened Included
arXiv 389 258 173
ACL Anthology 47 35 21
ACM Digital Library 26 15 5
IEEE Xplore 17 7 1
OpenReview 22 12 7
GitHub 7 4 2
Total 508 337 209
This search covered studies published between January 2020 and June 2025, ensuring inclusion of the most recent work on LLM robustness. Search strings were tailored to each database but generally combined terms related to “large language models”, “robustness”, “evaluation”, and “mitigation strategies”. A representative query structure was:
("large language model" OR "LLM" OR "foundation model")
AND (robustness OR reliable OR adversarial OR OOD OR evaluation OR mitigating robustness)
Figure 2: PRISMA flow diagram of the survey methodology
1.1.1 Inclusion Criteria
Studies were included if they met the following criteria:
  • Explicitly addressed robustness, reliability, or evaluation of LLMs.
  • Contained empirical experiments or benchmarks involving LLMs.
  • Published within the time range 2020–2025.
1.1.2 Exclusion Criteria
Studies were excluded if they:
  • Were duplicates or near-duplicate entries across sources or with similar concepts or results.
  • Did not contain experiments on LLMs (e.g., purely theoretical works, works on VLM).
  • Were published before 2020.
  • Were inaccessible or unretrievable.
Further, after completing this criterion, we applied backward snowballing (i.e., finding new publications by tracing references cited in this paper with some relevance). This process was repeated iteratively for each newly identified publication.
@@ 2 Robustness in LLMs
In the context of LLMs from various literatures, robustness refers to their ability to perform reliably and safely despite challenges like input variations, unexpected data, or adversarial attacks (which involve crafting specific inputs designed to mislead the model into producing incorrect or unintended outputs, despite the inputs appearing normal and being susceptible to humans). It is frequently discussed as a critical component within the broader framework of AI Trustworthiness. Trustworthy AI systems are expected to be reliable, safe, fair, transparent, accountable, and robust. Robustness underpins several of these qualities; for instance, a model prone to adversarial manipulation (lack of robustness) cannot be considered safe or reliable (wang2023robustnesschatgptadversarialoutofdistribution; aljohani2025comprehensivesurveytrustworthinesslarge). Similarly, a model that hallucinates or provides inconsistent answers under slightly different prompts lacks reliability, another facet closely tied to robustness (zhou2024trustworthinessretrievalaugmentedgenerationsystems). Research into the vulnerabilities of machine learning models, including LLMs, has often focused on adversarial robustness - the model’s resilience against inputs intentionally crafted to cause misclassification or undesired behaviour (zhao2025surveylargelanguagemodels).